the following information to answer questions 14 through 18. Assume Do-Nothing is an option. MARR-10% Alternative #1 Alternative #2 $205,000 $43,000 in year 1, First Cost $100,000 $35,000 per year increase of $3,000 per year S5,000 per years $10,000 per year Benefits O&M Costs Salvage Value for each of the following years $10,000 Useful Life ROR S15,000 8 years 10.85% 14. The NPW of Alternative #1 (necessary to make a decision between the two alternative) is closest to a) $3,249 $1,930 で) $543 e) $2,227 f)-$2,311 8) S723 h) $43I 1) $3,860 15. The increase in O&M costs that would result in a 10% ROR for Alternative #1 is closest to $483 per year 17 per year c)$813 per year d) $609 per year e) $407 per year f) $702 per year g)0&M costs need to decrease in order to earn a ROR of 10% h) not enough information to determine 6. The EUAB for Alternative #/2 is closest to... (Remember to treat the salvage value using the standard methodology a discussed in class.) a) $37,525 b) $43,811 c) $80,939 d) $72,373 e) $44,592 D$54,012 B) $55,323 17. The EUAC for Alternative #2 is closest to... (Remember to treat the salvage value using the standard methodology as discussed in class.) a) $48,4117 7,106 c) $50,782 d) $35,939 e) $23,714 #14
Answer #1

14. Ans: $1,930

Explanation:

NPW = -P -A(P/A, i, n) + A(P/A, i, n) + F(P/F, i, n)

         = -100,000 - 5,000(P/A, 10%, 4) + 35,000(P/A, 10%, 4) + 10,000(P/F, 10%, 4)

         = -100,000 - 5,000(3.170) + 35,000(3.170) + 10,000(0.6830)

         = -100,000 - 15,850 + 110,950 + 6,830

         = $1,930